OpenStudy (anonymous):

Find the surface area of each of the following cans rounding to the nearest whole number. a. Coffee can r= 7.6 cm h=16.3 b. soup can cm r=3.3 cm h=10 cm

OpenStudy (anonymous):

\[Surface Area = 2\pi r ^{2} + 2\pi r h \] A. \[SA = 2\pi (7.6)^{2} + 2\pi (7.6) (16.3) = 28.65 + 798.44 = 827.09cm\] B. \[SA = 2\pi (3.3)^{2} + 2\pi (3.3) (10) = 68.39 + 207.24 = 275.63cm\]

OpenStudy (anonymous):

I got for a. 1141cmsquared as the answer on the back of the book.

OpenStudy (anonymous):

The area for a cylinder is A:2pir^2+2pirh So for the 1st one we have a radius of 7.6 cm and height of 16.3> A:2pi(7.6^2)+2pi(7.6)(16.3)= 1,141.27777779cm^2<--- remember your units :) I'll let you figure out the next ;D

OpenStudy (anonymous):

Oh thanks JustHere!